OpenStudy (anonymous):

Given the equation:

OpenStudy (anonymous):

\[\sqrt{2x+1}=3\] x = 4, solution is extraneous x = 4, solution is not extraneous x = 5, solution is extraneous x = 5, solution is not extraneous

OpenStudy (anonymous):

\[\sqrt{x-5}+7=11\] x = 21, solution is extraneous x = 21, solution is not extraneous x = 81, solution is extraneous x = 81, solution is not extraneous

OpenStudy (anonymous):

\[-4\sqrt{x-3}=12\] x = 0, solution is not extraneous x = 0, solution is extraneous x = 12, solution is not extraneous x = 12, solution is extraneous

OpenStudy (anonymous):

I think the first one is x = 4, solution is not extraneous

OpenStudy (anonymous):

I think the second on is x = 81, solution is not extraneous

OpenStudy (anonymous):

and I have no idea about the third one

OpenStudy (usarmy3947):

so u need help with only the third one?

OpenStudy (anonymous):

Yeah...But am I right about one and two?

OpenStudy (anonymous):

@ganeshie8

OpenStudy (anonymous):

@madrockz

OpenStudy (anonymous):

I really need help

ganeshie8 (ganeshie8):

First question correct

ganeshie8 (ganeshie8):

Second question, INCORRECT> check again

OpenStudy (anonymous):

Yeah! So is it x = 81, solution is extraneous

ganeshie8 (ganeshie8):

how did u get x = 81 ?

OpenStudy (anonymous):

Umm...I think I Square rooted 9

OpenStudy (anonymous):

So I'm guessing its not 81

ganeshie8 (ganeshie8):

Second q :- \(\sqrt{x-5}+7=11\) put x = 21 see that it satisfies the equatio

ganeshie8 (ganeshie8):

thing inside sqrt becomes 21-5 = 16, so taking sqrt it becoems 4 4+7 = 11

ganeshie8 (ganeshie8):

so x = 21 is your solution, and its not extraneous. cuz it satisfied the given equaiton

OpenStudy (anonymous):

Wow...I was way off lol. Thanks...Do you understand how to do the third one?

ganeshie8 (ganeshie8):

:) Third question :- \(-4\sqrt{x-3}=12 \) divide both sides by -4 \( \sqrt{x-3}=3\) square both sides \( x-3 =9\) \(x = ?\)

OpenStudy (anonymous):

Would I add 3 then square...because if I just do 9 I get 81

ganeshie8 (ganeshie8):

add 3 both sides and stop there.

OpenStudy (anonymous):

oh so 12

ganeshie8 (ganeshie8):

x - 3 = 9 x = 12

ganeshie8 (ganeshie8):

yes, put x = 12 in given equation, and see if it satisfies or not

ganeshie8 (ganeshie8):

tbh, i wont do that cuz i see that sqrt can never become negative. but the given equation requires sqrt to be negative. so the solution x=12 in extraneous. it wont satisfy given equation

ganeshie8 (ganeshie8):

see if that makes some sense

OpenStudy (anonymous):

Yes defiantly so x = 12, solution is not extraneous

ganeshie8 (ganeshie8):

Careful

ganeshie8 (ganeshie8):

solution IS extraneous

OpenStudy (anonymous):

Oh wrong one....that's the part that keeps messing me up

OpenStudy (anonymous):

x = 12, solution is extraneous

ganeshie8 (ganeshie8):

yes

ganeshie8 (ganeshie8):

IF the solution DOES NOT satisfy given equation, we call it EXTRANEOUS

OpenStudy (anonymous):

Okay...I'll remember that! Thanks for helping! ^-^

ganeshie8 (ganeshie8):

np :)